The average radius of a molecule of lysozyme, an enzyme in tears, is 1.430 * 10−6 mm. What is its radius in micrometers (μm)? (Enter your answer in scientific notation.)
Answer #1

Explanation:

Step for unit conversion:

  1. Write down your problem.
  2. Find the conversion for required unit.
  3. Multiply your number by the conversion fraction.
  4. Cancel out your units.

Step 1: Analysis of unit

The given problem is

1.430 × 10^-6 mm

we know, here mm = millimeter = 10^-3 m or in simple word 1 mm =10^-3 m

we need to convert into um = micrometer =10^-6 m or 1um = 10^-6 m

Step 2: conversion of units

The problem is

1.430 × 10^-6 mm = 1.430 × 10^-6 mm × (10^-3 m / mm ) ×( um / 10^-6 m) = 1.430 × 10^-3 um

[ here, mm gets cancel out as well as m cancel out ]

Answer will be 1.430 × 10^-3 um